# Advancing Brainwave-Based Biometrics: A Large-Scale, Multi-Session Evaluation

### v1

@@ 3.3. Feature Extraction
To evaluate the effectiveness of metric learning loss functions in brainwave-based authentication, we used both softmax and pairwise approaches. Although PEERS is significantly larger than all previous brainwave datasets, it is still considerably smaller than face recognition datasets, such as MS-Celeb-1M, which contains over 10 million images from 100,000 identities (Guo et al., 2016). We selected SoftTripleLoss (Qian et al., 2019) and ArcFaceLoss (Deng et al., 2019) as our softmax-based approaches. SoftTripleLoss was chosen for its ability to handle intra-class variations by introducing multiple learnable centers per class, making it particularly effective for data with diverse feature distributions. On the other hand, ArcFaceLoss was selected for its angular margin constraints, which enforce strict inter-class separability and robustness against noise. We selected three pairwise-based loss functions: (1) Triplet Loss (Schroff et al., 2015), which optimizes embeddings by minimizing the distance between an anchor and a positive sample while maximizing the distance from a negative sample, serving as a widely used baseline; (2) LiftedStructureLoss (Oh Song et al., 2016), which leverages all positive and negative pairs within a batch simultaneously, penalizing violations of distance constraints and ensuring stable optimization; and (3) SupConLoss (Supervised Contrastive Loss) (Khosla et al., 2020), a recent and promising method that extends contrastive learning by explicitly using class labels to group all embeddings of the same class closer together while separating embeddings of different classes. This supervision allows SupConLoss to align embeddings more effectively across the entire batch, making it particularly suited for learning robust and discriminative feature representations.
3.4. Feature Comparison
After transferring and reducing brainwave samples from higher to lower dimensions via feature extraction, the next step involves comparing the extracted features and calculating a similarity score to determine verification attempts. Two main approaches are commonly used for this task. The first approach employs distance metrics, and we specifically select cosine similarity and Euclidean distance as they are widely used in biometric studies (Fallahi et al., 2024a; Jiang et al., 2023) and serve as distance functions in pairwise metric learning loss functions. The second approach involves training a binary classifier to distinguish between positive and negative samples for each subject. Positive samples correspond to data from the enrollment session, while negative samples are data from other subjects. However, negative samples used during evaluation must not overlap with those used in training, as this might lead to overestimation of performance by introducing familiarity with the subjects present in the training data—an unrealistic scenario in real-world authentication (Fallahi et al., 2024a, 2023). We select Support Vector Machines (SVM) and Random Forests, as they demonstrate superior performance in prior works (Arias-Cabarcos et al., 2023; Chaurasia et al., 2024), and are widely applied in brainwave authentication studies (Hu et al., 2021; Nagarajan et al., 2023; Shams et al., 2022).
3.5. Threat model
To evaluate the security of brainwave authentication systems as unlocking mechanisms for standalone devices (e.g., PCs, XR devices), we adopt zero-effort attacks (Mansfield and Wayman, 2002) as the primary adversary model, aligning with established biometric authentication standards (National Institute of Standards and Technology, 2023). A zero-effort attack in biometric authentication occurs when an adversary uses their own biometric sample in an attempt to gain unauthorized access to another individual’s account. In this context, the security objective of the biometric system is to minimize the probability that the authentication algorithm incorrectly verifies the adversary’s identity as the legitimate user. Specifically, the likelihood of successful verification of a falsely claimed identity, based on the adversary’s sensor readings, should be negligible. An active impersonation is not feasible at this stage. Because an attacker cannot infer any EEG-related information by observing the victim, nor is it feasible to collect EEG data from the victim without their consent. Even if EEG data were accessed, replay attacks can be prevented by implementing task-dependent brainwave authentication (Lin et al., 2018).
3.6. Performance Metrics
@@
We used the Equal Error Rate (EER) as a summary metric, indicating the point where the False Acceptance Rate (FAR) and the False Rejection Rate (FRR) are equal. The FAR represents the proportion of unauthorized attempts incorrectly accepted by the system, while the FRR refers to the proportion of legitimate attempts incorrectly rejected. EER has been widely used in biometric authentication papers (Rui and Yan, 2018; Ayeswarya and Singh, 2024) to report results and ensure comparability. However, in practical usage, it is preferred to keep FAR at very low rates to ensure system security, even at the cost of a higher FRR. Therefore, we also report the FRR at a FAR threshold of 1% for comparison with other brainwave authentication studies (Arias-Cabarcos et al., 2023; Fallahi et al., 2023; Navaratnam-Blair et al., 2022), as well as at 0.1% [^5] https://pages.nist.gov/800-63-3/sp800-63b.html and 0.01% [^6] https://pages.nist.gov/800-63-4/sp800-63b.html to align with biometric standards.
Table 1. Heatmap of benchmarking EER values for different combinations of feature extraction (PSD, PCA, SupConLoss, ArcFaceLoss) and comparison methods (RF, SVM, Cosine, Euclidean) in brainwave-based authentication, evaluated on 50 test subjects. NTS denotes the number of training and validation subjects used for training deep learning-based feature extractors.
NTS Feature Extractor Output Dimension Comparison Metric
Classifiers Distance Metrics
RF SVM Cosine Euclidean
0 Random Data 93 × 500 49.85 $\pm$ 0.22 49.99 $\pm$ 0.72 49.94 $\pm$ 0.88 49.92 $\pm$ 0.82
Raw EEG 93 × 500 39.49 $\pm$ 9.27 50.07 $\pm$ 0.71 38.07 $\pm$ 10.77 50.45 $\pm$ 13.61
PSD 93 × 129 32.53 $\pm$ 13.97 32.82 $\pm$ 12.23 39.30 $\pm$ 13.28 50.26 $\pm$ 12.28
PSD + Bins 93 × 5 30.54 $\pm$ 12.26 27.30 $\pm$ 14.33 37.52 $\pm$ 13.80 49.55 $\pm$ 12.45
PCA 93 × 5 37.51 $\pm$ 9.72 49.99 $\pm$ 0.86 37.03 $\pm$ 11.46 50.18 $\pm$ 13.15
PSD + PCA 93 × 5 29.16 $\pm$ 14.03 32.03 $\pm$ 12.13 45.47 $\pm$ 13.85 44.14 $\pm$ 15.52
245 Subjects PSD + SoftTripleLoss 128 30.95 $\pm$ 10.61 16.91 $\pm$ 10.52 15.54 $\pm$ 7.81 31.91 $\pm$ 15.07
PSD + ArcFaceLoss 128 32.83 $\pm$ 9.46 17.58 $\pm$ 10.48 18.52 $\pm$ 8.80 36.92 $\pm$ 16.89
PSD + Triplet Loss 128 35.25 $\pm$ 7.98 14.99 $\pm$ 7.57 15.28 $\pm$ 7.82 17.07 $\pm$ 8.38
PSD + LiftedStructureLoss 128 28.11 $\pm$ 11.33 12.67 $\pm$ 7.37 12.19 $\pm$ 7.15 17.34 $\pm$ 9.86
PSD + SupConLoss 128 27.73 $\pm$ 11.20 13.89 $\pm$ 8.36 12.34 $\pm$ 7.83 16.63 $\pm$ 9.58
SoftTripleLoss 128 29.88 $\pm$ 10.41 16.68 $\pm$ 10.29 15.59 $\pm$ 8.48 20.83 $\pm$ 12.18
ArcFaceLoss 128 27.53 $\pm$ 10.93 14.20 $\pm$ 8.34 14.22 $\pm$ 8.20 18.35 $\pm$ 10.96
Triplet Loss 128 35.61 $\pm$ 8.77 13.82 $\pm$ 8.78 13.04 $\pm$ 7.71 14.61 $\pm$ 9.28
LiftedStructureLoss 128 28.97 $\pm$ 11.59 12.41 $\pm$ 8.73 11.38 $\pm$ 7.24 14.25 $\pm$ 11.22
SupConLoss 128 25.62 $\pm$ 12.30 10.91 $\pm$ 6.84 10.75 $\pm$ 7.06 13.38 $\pm$ 9.85
4. Results and Discussion
@@
In order to understand multisession performance, we first require a pipeline that includes a feature extractor and a comparison metric, which we select through rigorous benchmarking. We then investigate important parameters such as test size, time interval, recording hardware, and subject embeddings visualized as distinct clusters through t-SNE plots. Next, to move toward consumer-grade devices, we analyze performance under a reduced number of channels and further contextualize our findings with current state-of-the-art methods and authentication standards to identify potential directions for future research.
4.1. Benchmarking Feature Extraction and Compression
@@
We have introduced the dataset and preprocessing methods; however, selecting suitable feature extraction and compression techniques is essential to finalize the authentication pipeline.
Experiment setup: To perform the benchmark, we need subject-exclusive sets for training and validation of the feature extractor, negative subjects for binary classification feature comparison, and test subjects for evaluation. The deep learning-based feature extraction approach was trained using 245 subjects, with 230 allocated for training and 15 for validation. Meanwhile, 50 subjects were specifically selected as negative samples for the shallow classifier feature comparison. Finally, we randomly selected 50 subjects to serve as the final test set.
Feature Extraction methods: Table 1 presents the results of the benchmark, beginning with the creation of random samples via a uniform distribution matching the size of the original data. This step ensured the absence of biases or misimplementations that could favor lower error results. The random samples achieved an expected baseline EER of approximately 50%. Classic feature extraction methods showed an improvement over this baseline but still yielded relatively high error rates. Among them, PSD features divided into bins demonstrated the best performance with a 27% EER. Transitioning to deep learning-based feature extraction methods resulted in significant performance improvements of 10% points of the EER or more. The best results were achieved by deep learning approaches utilizing raw data instead of preprocessed PSD features. Within these approaches, pair-based methods outperformed softmax-based ones. This advantage was expected due to the relatively larger dataset size compared to previous EEG authentication datasets, which allowed the methods to outperform PSD features. However, the dataset size remains smaller than the extensive datasets used for face detection, which often include thousands of subjects. Consequently, pair-based methods effectively leveraged the available data for better performance.
Comparison Methods: Table 1 demonstrates that SVM and cosine similarity outperform Random Forest and Euclidean distance as comparison functions. However, among the best results, which belong to Supervised Contrastive and Lifted Structure Loss feature extraction methods with raw and PSD input, the cosine similarity function consistently performs slightly better than SVM. The better performance with cosine similarity over SVM arises because deep learning loss functions optimize embeddings using distance metrics like cosine similarity or Euclidean distance in training stage.
Best Performing Authentication Pipeline Selection: The SupConLoss achieved the best results with an error rate of 10.75%, compared to 27.30% for the PSD approach, representing a 16.55% point improvement in the EER. Therefore, we selected SupConLoss as the feature extraction method for its superior performance. For feature comparison functions, cosine similarity, which does not require any prior data, outperformed the binary classification approach that requires 50 subjects as negative samples. Therefore, we selected cosine similarity for feature comparison in the remainder of the paper to apply within further investigations.
Figure 1. Number of sessions available in the test dataset after the enrollment session. Here, $W$ represents a week, $M$ represents a month, and $Y$ represents a year. The value following each symbol indicates the time period it represents, starting at a day after the end of the previous time unit and spanning the duration defined by the symbol ($W$, $M$, $Y$). For example, $W3$ refers to the number of sessions in the third week, beginning at the end of the second week (15th day) and ending on the 21st day. Similarly, $M2$ refers to the number of sessions during the second month, starting at the end of the first month (approximately the 31st day) and covering the next month. Note that $M\approx 30.42$ days and $Y=365.25$ days.
@@
Figure 2. EER vs. Number of Subjects: Each blue dot represents the average of 50 instances of randomly selecting $N$ subjects and calculating the EER. The dot size represents the standard deviation of the EER values across these 50 calculations. The trendline, calculated using ordinary least squares (OLS) regression, minimizes the squared differences between observed data points and the fitted line to reveal any trend between the number of subjects in test set and the average EER.
4.2. Detailed Performance Analysis
@@
With the best-performing pipeline identified, the EER remains significantly lower than the 50% chance level but still reaches 10.75%. This could indicate that brainwaves are not sufficiently unique for authentication or that other underlying issues need to be understood to develop solutions. Therefore, we aim to investigate which parameters influence multi-session authentication more deeply by addressing the following questions:
  1. (1)
@@
    How does altering the number of participants in an EEG-based authentication test set affect the overall comparability of the evaluation? To investigate this, we analyze how changing the number of test subjects impacts the EER in different scenarios.
  2. (2)
@@
Increasing the amount of test data provides deeper insights and allows for a more comprehensive evaluation of the multi-session authentication system. In the previous step, we selected cosine similarity as the feature comparison function. Consequently, the 50 subjects previously designated as negative samples for binary classification can now be incorporated into our evaluation, as they were not utilized in any part of the current pipeline. This inclusion allows us to expand the test set to 100 subjects, enabling a more robust assessment of system performance. We observe that using SupConLoss and cosine similarity results in an EER of 10.78% ± 6.67% on the 100 test subjects, showing only a marginal 0.03% increase relative to the 50-subject test set. We could also use the data from these 50 subjects to improve the feature extractor, which would likely result in a slightly lower EER. However, this approach would limit our ability to investigate scenarios with fewer sessions across different time intervals or a smaller number of test subjects per headset.
@@
Figure 3. Relationship between the time interval between sessions and EER. $D$ represents a day, $W$ a week, $M$ a month, and $Y$ a year. The value after each symbol indicates the start time interval (except for days). For example, $W1$ represents sessions with a time interval of 8–14 days after the first session.
4.2.1. Effect of the Number of Test Subjects on EER:

### v2

@@ III-C Feature Extraction
For softmax-based approaches, we selected two loss functions: (1) SoftTripleLoss, which addresses intra-class variations by introducing multiple learnable centers per class; and (2) ArcFaceLoss, which applies angular margin constraints to improve inter-class separability and robustness to noise. For distance-based approaches, we selected three loss functions: (1) Triplet Loss [40], which minimizes the distance between an anchor and a positive sample while maximizing the distance from a negative sample; (2) LiftedStructureLoss [41], which uses all positive and negative pairs within a batch and penalizes violations of distance constraints; and (3) SupConLoss (Supervised Contrastive Loss) [42], which extends contrastive learning by using class labels to group embeddings of the same class together and separate embeddings of different classes. This supervision helps SupConLoss align embeddings more effectively across the batch.
For each loss function, different encoder architectures were explored to optimize feature extraction. Specifically, we implemented six backbone EEG models — ResNet1D [43], ShallowNet [44], DeepConvNet [44], EEGNet [45], LSTM [46], and GRU [46] — as encoder networks. These models were selected to cover a diverse range of architectures, including convolutional and recurrent networks, and to balance model complexity and representational capacity for EEG signal processing.
III-D Feature Comparison
After reducing brainwave samples to lower dimensional embeddings through feature extraction, the extracted features should be compared, and a similarity score calculated to evaluate verification attempts. Two main approaches can be used. The first approach applies distance metrics; we select Euclidean, Cosine, and Manhattan distances because Euclidean captures absolute differences, Cosine measures angular similarity independent of magnitude, and Manhattan is more robust to outliers [4, 47]. The second approach trains binary and one-class classifiers to distinguish between positive and negative samples for each subject. However, binary classifiers require negative samples, which should not come from subjects included in the evaluation. We employ several binary classifiers, including logistic regression, linear discriminant analysis (LDA), support vector machine (SVM) with linear kernel, stochastic gradient descent (SGD) classifier with radial basis function (RBF) kernel approximation, and random forests. One-class SVM is also tested, although it is less suitable when negative classes are available; it is included due to its use in related work [12]. As shown in face recognition, distance-based approaches are expected to perform better if sufficient data is available to train the feature extractor [40], since metric learning aims to minimize distances between similar samples. The most suitable comparison method will be selected via systematic hyperparameter tuning, as described in the Section IV-A.
III-E Threat model
Since EEG signals are not externally observable, shoulder surfing is not possible, and the adversary cannot impersonate a victim. Accordingly, authentication studies typically assume that adversaries attempt to gain access using their own brainwave recordings [8, 5]. Meanwhile, our goal is to investigate the reliability of brainwave-based authentication over time (multi-session setting), and to identify potential bottlenecks affecting this reliability. The low reliability of brainwave signals across sessions contributes to a high false acceptance rate, which poses a significant security threat and can be assessed using zero-effort attacks [48]. A zero-effort attack in biometric authentication occurs when an adversary uses their own biometric sample in an attempt to gain unauthorized access to another individual’s account. In this context, the security objective of the biometric system is to minimize the probability that the authentication algorithm incorrectly verifies the adversary’s identity as the legitimate user. Specifically, the likelihood of successful verification of a falsely claimed identity, based on the adversary’s sensor readings, should be negligible. Therefore, zero-effort attacks are commonly used in studies on multi-session reliability involving brainwave data [12, 16, 10, 11, 14, 13].
III-F Performance Metrics
@@
We used the Equal Error Rate (EER) as a summary metric, indicating the point where the False Acceptance Rate (FAR) and the False Rejection Rate (FRR) are equal. The FAR represents the proportion of unauthorized attempts incorrectly accepted by the system, while the FRR refers to the proportion of legitimate attempts incorrectly rejected. EER has been widely used in biometric authentication papers [25, 49] to report results and ensure comparability. However, in practical usage, it is preferred to keep FAR at very low rates to ensure system security, even at the cost of a higher FRR. Therefore, we also report the FRR at a FAR threshold of 1% for comparison with other brainwave authentication studies [5, 9, 50], as well as at 0.1% [^5] https://pages.nist.gov/800-63-3/sp800-63b.html and 0.01% [^6] https://pages.nist.gov/800-63-4/sp800-63b.html to align with biometric standards.
TABLE II: Results of top 1,000 hyperparameter trials on validation set (15 subjects): pipeline counts, average EER, and best EER among all trials.
Pipeline Count Avg Best
Feature Extraction
ResNet1D + SupConLoss 98 10.20 8.05
ResNet1D + ArcFaceLoss 65 10.65 8.09
ResNet1D + LiftedStructureLoss 107 10.78 8.39
ShallowNet + SoftTripletLoss 77 11.13 8.97
ResNet1D + TripletMarginLoss 99 11.36 9.20
ShallowNet + LiftedStructureLoss 34 10.85 9.25
ShallowNet + SupConLoss 77 11.50 9.38
ResNet1D + SoftTripletLoss 54 11.79 9.66
ShallowNet + ArcFaceLoss 72 12.12 9.69
DeepConvNet + SoftTripletLoss 90 12.19 10.23
GRU + SoftTripletLoss 36 11.43 10.35
LSTM + SoftTripletLoss 28 11.87 10.72
DeepConvNet + ArcFaceLoss 26 12.10 11.11
ShallowNet + TripletMarginLoss 47 12.30 11.50
GRU + SupConLoss 28 12.62 12.04
DeepConvNet + SupConLoss 10 12.69 12.34
LSTM + LiftedStructureLoss 9 12.71 12.59
DeepConvNet + LiftedStructureLoss 7 12.68 12.63
GRU + LiftedStructureLoss 8 12.69 12.65
EEGNet + LiftedStructureLoss 9 12.93 12.75
GRU + ArcFaceLoss 19 12.89 12.79
EEGNet + ArcFaceLoss 0 0.00 13.18
LSTM + ArcFaceLoss 0 0.00 13.21
EEGNet + SupConLoss 0 0.00 13.36
LSTM + SupConLoss 0 0.00 13.61
DeepConvNet + TripletMarginLoss 0 0.00 13.88
EEGNet + SoftTripletLoss 0 0.00 14.49
EEGNet + TripletMarginLoss 0 0.00 15.00
LSTM + TripletMarginLoss 0 0.00 15.55
GRU + TripletMarginLoss 0 0.00 18.74
PSD + AR 0 0.00 19.23
AR 0 0.00 19.43
PSD 0 0.00 29.34
Feature Comparison
Euclidean Distance 86 11.18 8.05
Cosine Distance 168 10.75 8.05
Manhattan Distance 105 10.80 8.17
Logistic Regression 171 11.70 8.36
LDA 222 11.67 8.56
Linear SVM 106 11.95 8.86
SGD RBF Kernel 80 11.74 9.58
Random Forest 62 12.45 11.49
One-Class SVM 0 0.00 21.68
IV Results and Discussion
@@
In order to evaluate brainwave authentication performance over time, we first select a pipeline comprising a feature extractor and a comparison metric through hyperparameter tuning. We then investigate important parameters such as test size, time interval, recording hardware, and subject embeddings visualized as distinct clusters through t-SNE plots. Next, to move toward consumer-grade devices, we analyze performance with a reduced number of channels and contextualize our findings with state-of-the-art methods and authentication standards to identify potential directions for future research.
IV-A Benchmarking Feature Extraction and Compression
@@
After introducing dataset and preprocessing methods we now want to turn to selecting suitable feature extraction and compression techniques, that are essential to complete the authentication pipeline.
Experiment setup: Our claim is that the proposed method supports both inter-subject and inter-session generalization. Therefore, this is reflected by the data split and evaluation protocol. We divided the 345 available subjects into non-overlapping sets: 230 subjects for training the deep learning-based feature extractor, 15 for validation, and 100 randomly selected subjects for evaluation, on whom system performance is measured. In all experiments, unless specified otherwise, the first session of each subject was used for enrollment, and the remaining sessions of the same subject for verification. This setup ensures inter-subject evaluation and shows that the model does not require retraining when new users are added. Additionally, separating enrollment and verification by session enables the assessment of inter-session generalization.
Hyperparameter tuning: From each combination of model architectures and loss functions (6×5) mentioned in the methods (Section III), we trained a model using hyperparameter tuning on the optimizer, learning rate, batch size, and embedding dimension based on the training set, guided by performance on the validation set. The entire process took approximately 144 hours on four NVIDIA GeForce RTX 4090 GPUs.
After obtaining the feature extraction models, we conducted general hyperparameter tuning to select the best combination of feature extraction and feature comparison methods. The tuning process was guided by the validation set and employed a random search over 10,000 trials. To provide a clearer overview, Table II reports the results for the top 10% of trials on the validation set, thereby reducing the influence of high-error outliers on the average results. The hyperparameter tuning identified the “ResNet1D + SupConLoss + Euclidean distance” pipeline as the best configuration. This pipeline was then evaluated on a test set of 100 subjects, resulting in an EER of 10.28% ± 6.3%. For the remainder of this paper, we use this pipeline as the default authentication pipeline.
Feature Extraction Methods: Table II shows that our dataset is sufficient for deep learning approaches to outperform handcrafted features, which result in significantly higher error. Among the encoders, ResNet1D consistently achieved the top three ranks, followed by ShallowNet, which outperformed the remaining models. This suggests that convolutional networks generally perform better than recurrent architectures. Among the convolutional models, simpler architectures like ShallowNet may benefit from reduced overfitting, while ResNet1D leverages residual connections to achieve better performance than models such as DeepConvNet. Although EEGNet is relatively shallow, its specialized filters for frequency and spatial information did not translate into strong performance in our evaluation. Regarding loss functions, although SoftTripleLoss—with its multiple learnable centers per class—performed best with shallower and recurrent backbones (ShallowNet, DeepConvNet, LSTM, GRU), ResNet1D combined with SupConLoss achieved the highest overall performance across all architectures. SupConLoss combines the strengths of distance-based and softmax-based approaches by leveraging all positive and negative pairs in the batch, guided by class labels. This batch-level supervision enforces strong intra-class cohesion and inter-class separation, which ResNet1D—with its higher capacity—can fully exploit to learn highly discriminative EEG embeddings
Comparison Methods: The results (Table II) suggest that, in general, distance-based functions outperform shallow classifier approaches for feature comparison. This is not surprising, as loss functions in metric learning are designed to minimize distances between embeddings, typically using cosine or Euclidean distance. In contrast, shallow classifiers either require negative samples for training or, in the case of one-class classification, often show relatively poor performance due to the instability associated with one-class models. This is also why state-of-the-art face recognition methods use distance functions for feature comparison.
Figure 1: Number of sessions in the evaluation dataset post-enrollment. $W$, $M$, and $Y$ denote week, month, and year. Labels (e.g., $W3$) mark the number of sessions in that time unit, starting the day after the previous unit ends (e.g., $W3$: days 15–21).
@@
Figure 2: EER vs. Number of Subjects: Each blue dot represents the average of 50 instances of randomly selecting $N$ subjects and calculating the EER. The dot size represents the standard deviation of the EER values across these 50 calculations.
IV-B Detailed Performance Analysis
@@
With the best-performing pipeline identified, the EER remains significantly lower than the 50% chance level but still reaches 10.28%, which is higher than the performance required for real-world deployment. This result may indicate that brainwaves are not sufficiently unique for authentication or that other underlying factors need to be understood to develop effective solutions. Therefore, we aim to investigate which parameters influence multi-session authentication more deeply by addressing the following research questions:
  1. 1.
@@
    How does altering the number of participants in an EEG-based authentication evaluation set affect the overall comparability of the evaluation? To investigate this, we analyze how varying the number of evaluation subjects from 2 to 100 affects the EER.
  2. 2.
@@
Figure 3: Relationship between the time interval between sessions and EER. $D$ represents a day, $W$ a week, $M$ a month, and $Y$ a year. The value after each symbol indicates the start time interval (except for days). For example, $W1$ represents sessions with a time interval of 8–14 days after the first session.
IV-B1 Effect of the Number of Evaluation Subjects on EER